# A Guide to Stochastic Optimisation for Large-Scale Inverse Problems

### v1

@@ 4.2   Hyperparameter Tuning
Parameter tuning has been an important yet long-standing challenge in stochastic optimisation. Due to their random nature, most algorithmic parameters are usually much harder to be either selected a-priori or tuned on-the-fly, comparing to their deterministic counterparts. The typical algorithmic parameters which commonly need to be tuned include:
@@
Step-size selection is an important problem in stochastic optimisation, and has been a driving force behind the development of advanced stochastic algorithms, e.g. variance-reduced stochastic algorithms. On the one hand, stochastic algorithms inherit the step-size issues that deterministic algorithms exhibit, that is it requires knowing the Lipschitz constant, which may not be available, such as too expensive evaluate or highly spatially varying. On the other hand, step-sizes with theoretical guarantees are often too conservative. While in practice, step-sizes exceed theoretical limit often work and provide even faster performance. For example, in the first experiment of Section 5, step-size $1/(1.75nL_{\max})$ is used for SAGA under which the algorithm converges. Similarly, primal-dual based algorithms also require proper balancing of the primal and dual step-sizes to yield good performance. Overall, the step-size selection for stochastic optimisation algorithms needs more dedicated research.
@@
While the line-search schemes can usually be implemented for deterministic gradient methods such as PGD or FISTA [BT09], applying these schemes for stochastic gradient methods are difficult in general both in theory and in practice. Traditional line search methods require evaluating the full objective is undesirable since it requires using all the data. Using the corresponding stochastic variants can perform well in practice but lacks theoretical guarantees [GSBR20]. Recently, a number breakthroughs have been made for on-the-fly tuning for the step-size of stochastic gradient methods applied in machine learning [DM23], but not yet fully adapted to the proximal setting which is prevalent in inverse problems.
@@
Selecting the subset pattern is also essential for optimally applying stochastic gradient methods. As discussed in [TEGD20], the subset pattern influences the gradient Lipschitz constants on subset, and hence change the admissible step-size and convergence speed. The influence of the subset selection has also been discussed in [GDE22] for parallel MRI. The choice of subset pattern is reasonably well-understood for some inverse problems like full-angle 2D CT. However, for many other problems like limited-angle CT, 3D PET etc, this problem has not been resolved.
@@
Accelerated stochastic gradient methods do not automatically adapt to the strong-convexity or hidden restricted strong-convexity which is prevalent when using sparsity-inducing regularizers. Without restarting, these methods would still convergence fast but only initially. When approaching high-accuracy, the algorithms start to oscillate around the solution and which slows down convergence. To address these, adaptive restart schemes have been developed to accelerate momentum-based methods [TGBD18, FQ20].
@@
In practice, variance-reduced methods such as SVRG and SAGA, are slower than vanilla SGD at the start. Replacing the initial iterations of variance-reduced methods with a warm-start phase running vanilla SGD, a faster convergence at the beginning of iterations has been observed [TAK+23, KR17]. However, it is unclear how long this phase should be a-priori.
@@ 4.3   Memory Considerations
Vanilla SGD does not require storing any additional pieces of information, and thus has the same memory footprint as GD, $\mathcal{O}(|\mathcal{X}|)$ required to store the current iterate.
SAG and SAGA, in their standard form, require storing the table of computed gradients. The sum of the elements in the table then either needs to be stored or computed anew in each iteration. Thus, the memory overhead of the standard implementation of SAG and SAGA is $\mathcal{O}(n|\mathcal{X}|)$, which can be prohibitively large if $n$ and $|\mathcal{X}|$ are large. In particular, in large scale CT the dimensionality $|\mathcal{X}|$ can be very large. However, memory requirements can be significantly lowered in certain mitigating scenarios. Consider linear models [DBLJ14, GSBR20], i.e. when $h_{i}=\overline{h}_{i}\circ A_{i}$ for all $i=1,\ldots,n$ and some linear operators $A_{i}$. This is for example the case for the least-squares data fidelity or the Kullback–Leibler divergence (see Section 1). The gradient of $h_{i}$ then satisfies $\nabla h_{i}(x)=A_{i}^{\ast}\nabla\overline{h}_{i}(A_{i}x)$. In other words, instead of storing $\nabla h_{i}(x)$ we can store $\nabla\overline{h}_{i}(A_{i}x)$. As the original SAGA, each iteration the application of $A_{i}$ and $A_{i}^{*}$, so no extra computational cost is introduced. The modified SAGA algorithm is written in Algorithm 12
@@ 7       $x^{(k+1)}=\operatorname{prox}_{\tau g}\big{(}x^{(k)}-\tau\tilde{\nabla}h(x^{(k)})\big{)}$
The standard implementation of SVRG (see Algorithm 9) requires storing the reference iterate $\overline{x}^{(j)}$ and the full gradient $\overline{\gamma}^{(j)}$, resulting in a memory of $\mathcal{O}(|\mathcal{X}|$) which is independent of the number of subsets and comparable to GD. However, this comes at the cost of computing two stochastic gradients per iteration, one at the previous iterate and another at the reference iterate. In some applications, the memory overhead of storing $\nabla h_{i}(\overline{x}^{(j)})$ might be preferable over the extra gradient computation $\nabla h_{i}(\overline{x}^{(j)})$. As for SAG/SAGA, we can derive a modified version of SVRG which has the same memory footprint $\mathcal{O}(|\mathcal{X}|+|\mathcal{Y}|)$ and the same computational cost as the modified SAGA algorithm.
@@ 5   Numerical Examples
In this section we present a brief numerical study, showcasing representative stochastic optimisation algorithms discussed in this review, and comparing them against their deterministic counterparts. Note that this numerical study is not intended to provide a comprehensive and thorough comparison of algorithmic performance. Instead, our aim is to show that stochastic algorithms can indeed be beneficial for certain problems and illustrate their behaviour. The code to reproduce all results in this paper will be made available upon acceptance of this manuscript.
@@ 5.1   Sparse Spikes Deblurring
In this first experiment, we consider the toy problem of deblurringw where the ground truth are sparse spikes and compare PGD and SAGA. The purpose of this example is to find possible factors affect the performances of the two algorithms.
@@
Let $u^{\dagger}\in\mathbb{R}^{d},d=200$ be a sparse vector, $K\in\mathbb{R}^{d\times d}$ be the matrix representation of a blur with uniform kernel of size $\kappa$ normalised such that $\|K\|=1$ and periodic boundary conditions. We consider noisy data $v=Ku+\eta$, where $\eta\in\mathbb{R}^{d}$ is random Gaussian noise. To recover $u^{\dagger}$ from the observation $v$, we minimise $\Phi(x)=g(x)+h(x)$ with $x=u$, $g(x)=\mu{|\kern-1.125pt|}x{|\kern-1.125pt|}_{1}$ and $h(x)=\frac{1}{2}{|\kern-1.125pt|}Kx-v{|\kern-1.125pt|}_{2}^{2}$. Written into the finite-sum form, we have $h(x)=\sum_{i=1}^{n}h_{i}(x)$ where $n=d$, $h_{i}(x):=\frac{1}{2}(K_{i}x-v_{i})^{2}$ and $K_{i}\in\mathbb{R}^{1\times n},i=1,...,n$ is the $i$-th row of $K$. Note that $L=1$, ${|\kern-1.125pt|}K_{i}{|\kern-1.125pt|}=1/\sqrt{\kappa},i=1,...,n$ and thus $L_{\max}=1/\kappa$.
We consider $100$ kernel sizes with $\kappa=2i-1$ for $i=1,\ldots,100$, and random Gaussian noise $\eta\sim\mathcal{N}(0,\sigma^{2}I)$, with $10$ different noise levels, taken at equidistant points in $[10^{-4},10^{-1}]$. The regularisation parameter $\mu$ is chosen as $\mu={|\kern-1.125pt|}K^{*}v{|\kern-1.125pt|}_{\infty}/3$. For PGD and SAGA, two step-sizes are chosen for each method: 1) for PGD, we choose $\tau=1/L$ (denoted PGD1) and $\tau=1.9/L$ (denoted PGD2); 2) for SAGA, we choose $\tau=1/(3nL_{\max})$ (denoted SAGA1) and $\tau=1/(1.75nL_{\max})$ (denoted SAGA2).
@@
Given a kernel size and noise levels we first compute a high-precision solution $x^{\ast}$ and then run PGD and SAGA using the stopping criterion ${{|\kern-1.125pt|}x^{(k)}-x^{\ast}{|\kern-1.125pt|}}/{{|\kern-1.125pt|}x^{\ast}{|\kern-1.125pt|}}\leq 10^{-4}.$ We record the number of data passes needed to reach that accuracy, denoted as $N_{\text{PGD}}$ and $N_{\text{SAGA}}$ for PGD and SAGA, respectively, and compute the ratio $N_{\text{SAGA}}/N_{\text{PGD}}$ averaged over 10 noise levels. Since for each method, we have two different choices of step-sizes, therefore in total we have $4$ ratios, which are provided in Figure 1, accompanied with an extra line $1001/\kappa_{i},i=1,\ldots,$ 100.
@@
From Figure 1, we obtain the following observations. First, when the kernel size is one, then $K$ is the identity matrix. For this case, PGD with $\tau=1/L$ converges in one step, hence the ratio $N_{\text{SAGA}}/N_{\text{PGD}}$ is very large. Also, SAGA1 is faster than SAGA2. One reason for this when for this mall kernel size, the gradients of $h_{i}$ are least correlated, and larger step-size leads to larger variance and slower performance. Second, for kernel sizes between $5$ and $199$, we observe that the ratio $N_{\text{SAGA}}/N_{\text{PGD}}$ is proportional to $1/\kappa$. Moreover, for both PGD and SAGA, larger step-size leads to faster performance. To see this, notice that the orange line $N_{\text{SAGA1}}/N_{\text{PGD2}}$ is above the red line $N_{\text{SAGA1}}/N_{\text{PGD1}}$, indicating that PGD1 needs more data pass to reach stopping criterion. While the green line below the red line means SAGA1 requires more data pass to reach the stopping criterion than SAGA2.
@@
Figure 1: The coloured lines represent the ratio of the number of data passes for SAGA and PGD, i.e. $N_{\text{SAGA}}/N_{\text{PGD}}$, for different step-size choices. See text for more details.
5.2   Parallel CT
@@
In this set of experiments we investigate tomographic reconstruction of a Shepp–Logan phantom of size $128\times 128$. The measurements follow the forward problem $v=Ku^{\dagger}+\eta$, where $\eta$ is a random sample from $\mathcal{N}(0,\sigma^{2}I)$, and the forward operator $K$ is defined by parallel CT with $240$ angles from the $[0,\pi)$ angle range.
@@
We compare the performance of a deterministic algorithm PGD (Algorithm 1) with SGD (Algorithm 7) and SAGA (Algorithm 8). As discussed in Section 3, when solving inverse problems it is not common to utilise fully stochastic approaches, in the sense that it is not common to use randomly constructed batches but rather to pre-partition measurement data into batches, and then one element of the partition is sampled at random in each iteration.
@@
PGD minimises $\Phi(x)=g(x)+h(x)$, where $x=u$, $g(x)=\lambda\text{TV}(x)$, and $h(x)=\frac{1}{2}\|Kx-v\|^{2}$. SGD and SAGA minimise $\Phi(x)=g(x)+\sum_{i=1}^{n}h_{i}(x)$, where $n\in\mathbb{N}$, and $h_{i}(x)=\frac{1}{2}\|K_{i}x-v_{i}\|^{2}$ is selected such that $\sum_{i=1}^{n}h_{i}(x)=h(x)$. This means, we define $K_{i}$, for $1\leq i\leq n$, as the CT operator that uses $240/n$ angles from the range $[i/n,\pi)$ by taking every $n$-th angle starting with $i/n$ giving a $43n/$ 4-degree angle separation. This is sometimes called the equidistant or staggered partition, and results in all the operator norms $\|K_{i}\|$ to be of a nearly identical value. Moreover, the number of batches is typically a divisor of the number of angles, which ensures that all the batches are of the same size.
@@
(c) Reconstructions after 10 data passes with $n=60$
@@
In Figure 2 we show the performance of these 3 algorithms in terms of the sub-optimality of the objective value: $\Phi(x^{(k)})-\Phi(x^{\ast})$, where $\Phi(x^{\ast})$ is a reference solution computed with FISTA. Figure 2 shows a significant speed up with SGD and SAGA over PGD. SGD is the fastest initially, but it then slows down after 5-10 data passes. On the other hand, SAGA is steady and fast throughout the optimisation trajectory and the increased variance it presents in initial iterations disappears already after around 20 data passes. Figure 2a and Figure 2b also investigate the effect of the number of subsets: the lightest shade shows a lower number of subsets ($n=10$), normal shade shows ($n=60$), and the darkest shade shows a high number of subsets ($n=240$). We observe that increasing the number of subsets has a dual effect: it increases the variance of the optimisation trajectory for stochastic algorithms, but it greatly increases the speed. In Figure 2c we show the reconstructions of the 4 studied algorithms obtained after $10$ data passes. The reconstruction with PGD presents as oversmooth, and fails to adequately reconstruct the finer image detail. For example, consider the zoomed in region where stochastic algorithms manage to recover two distinct objects (one of a lower and the other of a higher intensity), whereas PGD does not.
@@
In Figure 3 we compare the effect of the strategy for accessing the function indices. Specifically, we compare SGD and SAGA, with indices sampled uniformly at random (with replacement), and IGD and IAGA, which are incremental gradient modifications of SGD and IAGA that access indices according to a deterministic Herman–Meyer order [HM93]. The results show that the behaviour highly depends on the number $n$ of subsets and the algorithm. Comparing non-variance reduced algorithm reveals that the deterministic strategy can outperform the stochastic strategy, though the differences are not significant. However, for variance reduced algorithms the differences are significantly more severe. When $n$ is small, SAGA outperforms IAGA and presents a much smaller variance, though the variance with IAGA reduces significantly within the first $50$ data passes. When $n$ is large, SAGA is significantly better than IAGA.
5.3   Fan-Beam CT
In the last set of experiments we take the central slice of a cone beam CT scan of a walnut [Mea20]. Full sinogram data has 721 projections in the $[0,2\pi)$ range. To reduce the computational costs we consider only $360$ projections in the $[0,2\pi)$ range, with a $1^{\circ}$ angle separation. The reconstructed images are of size $280\times 280$.
Noisy measurements are generated by corrupting the sinogram (the measurement data) following the Beer–Lamberts law. We first compute the expected number of counts via $N_{I}=\text{Pois}(I_{0}\exp(-v))$, where $v$ is the loaded (noiseless) sinogram and $I_{0}$ the intensity of the initial beam. The noisy measurement is then computed by first correcting all 00 entries in $N_{I}$ to $1$, and then applying linearisation (post-log measurements) as $v^{\delta}=-\ln(N_{I}/I_{0})$. We consider three intensity levels: low intensity ($I_{0}=50$), medium intensity ($I_{9}=250$), and high intensity ($I_{0}=5000$), simulating high to low noise. The objective is defined as

### v2

@@ 4.2   Hyperparameter Tuning
Parameter tuning has been an important yet long-standing challenge in stochastic optimisation. Due to their random nature, most algorithmic parameters are usually much harder to be either selected a-priori or tuned on-the-fly, compared to their deterministic counterparts. The typical algorithmic parameters which commonly need to be tuned include:
@@
Step-size selection is an important problem in stochastic optimisation, and has been a driving force behind the development of advanced stochastic algorithms, e.g. variance-reduced stochastic algorithms. On the one hand, stochastic algorithms inherit the step-size issues that deterministic algorithms exhibit, such as requiring to know the Lipschitz constant, which may not be available, too expensive to compute accurately, or highly spatially varying. On the other hand, step-sizes with theoretical guarantees are often too conservative and in practice, step-sizes larger than the theoretical limit often work well and provide an even faster performance. For example, in the first experiment of Section 5, step-size $1/(1.75nL_{\max})$ is used for SAGA under which the algorithm converges. Similarly, primal-dual based algorithms require proper balancing of the primal and dual step-sizes to yield good performance. Overall, the step-size selection for stochastic optimisation algorithms needs more dedicated research.
@@
While the line-search schemes can usually be implemented for deterministic gradient methods such as PGD or FISTA [BT09], applying these schemes for stochastic gradient methods is in general difficult, both in theory and in practice. Traditional line search methods that require evaluating the full objective are undesirable, since this would require using all the data. Using the corresponding stochastic variants can perform well in practice but lacks theoretical guarantees [GSBR20]. Recently, a number of breakthroughs have been made for on-the-fly tuning of the step-size for stochastic gradient methods when applied in machine learning [DM23]. However, this has not yet been fully adapted to the proximal setting which is prevalent in inverse problems.
@@
Selecting the subset pattern, i.e. partitioning of the forward operator, is essential for optimally applying stochastic gradient methods. As discussed in [TEGD20], the subset pattern determines the gradient Lipschitz constants on subset, and hence affects the admissible step-size and convergence speed. The influence of the subset selection has also been discussed in [GDE22] for parallel MRI. The choice of a suitable subset pattern is reasonably well-understood for some inverse problems like full-angle 2D CT. However, for many other problems like limited-angle CT, 3D PET etc, this problem has not been resolved.
@@
Accelerated stochastic gradient methods do not automatically adapt to the strong-convexity or hidden restricted strong-convexity which is prevalent when using sparsity-inducing regularisers. Without restarting, these methods would still show fast convergence but only initially. When approaching high-accuracy solutions, the algorithms start to oscillate around the solution, slowing down the convergence. To address this, adaptive restart schemes have been developed to accelerate momentum-based methods [TGBD18, FQ20].
@@
In practice, variance-reduced methods such as SVRG and SAGA, are slower than vanilla SGD at the start. Replacing the initial iterations of variance-reduced methods with a warm-start phase by running vanilla SGD, a faster convergence at the beginning of iterations has been observed [TAK+23, KR17]. However, it is unclear how long this phase should be a-priori.
@@ 4.3   Memory Considerations
Vanilla SGD does not require storing any additional pieces of information, and thus has the same memory footprint as GD: $\mathcal{O}(|\mathcal{X}|)$ required to store the current iterate.
SAG and SAGA, in their standard form, require storing the table of computed gradients. The sum of the elements in the table then either needs to be stored or computed anew in each iteration. Thus, the memory overhead of the standard implementation of SAG and SAGA is $\mathcal{O}(n|\mathcal{X}|)$, which can be prohibitively large if $n$ and $|\mathcal{X}|$ are large. In particular, in large scale CT the dimensionality $|\mathcal{X}|$ can be very large. However, memory requirements can be significantly lowered in certain mitigating scenarios. Consider linear models [DBLJ14, GSBR20], i.e. when $h_{i}=\overline{h}_{i}\circ A_{i}$ for $i=1,\ldots,n$ and linear operators $A_{i}$. This is for example the case for the least-squares data fidelity or the Kullback–Leibler divergence (see Section 1). The gradient of $h_{i}$ then satisfies $\nabla h_{i}(x)=A_{i}^{\ast}\nabla\overline{h}_{i}(A_{i}x)$. In other words, instead of storing $\nabla h_{i}(x)$ we can store $\nabla\overline{h}_{i}(A_{i}x)$. As in the original SAGA, each iteration needs one application of $A_{i}$ and $A_{i}^{*}$, thus no extra computational cost is introduced. The modified SAGA algorithm is written in Algorithm 12
@@ 7       $x^{(k+1)}=\operatorname{prox}_{\tau g}\big{(}x^{(k)}-\tau\tilde{\nabla}h(x^{(k)})\big{)}$
The standard implementation of SVRG (see Algorithm 9) requires storing the reference iterate $\overline{x}^{(j)}$ and the full gradient $\overline{\gamma}^{(j)}$, resulting in a memory of $\mathcal{O}(|\mathcal{X}|$) which is independent of the number of subsets and comparable to GD. However, this comes at the cost of computing two stochastic gradients per iteration, one at the previous iterate and another at the reference iterate. In some applications, the memory overhead of storing $\nabla h_{i}(\overline{x}^{(j)})$ might be preferable over the extra gradient computation $\nabla h_{i}(\overline{x}^{(j)})$. As for SAG(A), we can derive a modified version of SVRG which has the same memory footprint $\mathcal{O}(|\mathcal{X}|+|\mathcal{Y}|)$ and the same computational cost as the modified SAGA.
@@ 5   Numerical Examples
In this section we present a brief numerical study, showcasing representative stochastic optimisation algorithms discussed in this review, and comparing them against their deterministic counterparts. Note that this numerical study is not intended to provide a comprehensive and thorough comparison of algorithmic performance. Instead, our aim is to show that stochastic algorithms can indeed be beneficial for certain problems and to illustrate their behaviour. The code to reproduce all results in this paper will be made available upon acceptance of this manuscript.
@@ 5.1   Sparse Spikes Deblurring
We consider a toy deblurring problem where the ground truth are sparse spikes and the forward operator uses a uniform blurring kernel. The goal of this example is to compare PGD and SAGA and investigate how is the performance of those two algorithms affected by the size of the blurring kernel and step-size regime. The numerical results were produced with MATLAB.
@@
Let $u^{\dagger}\in\mathbb{R}^{d},\ d=1,000$ be a sparse vector with $20$ evenly distributed non-zeros entries, representing the ground truth, $K\in\mathbb{R}^{d\times d}$ be the matrix representation of a blur with uniform kernel of size $\kappa$, with periodic boundary conditions, which is normalised such that $\|K\|=1$. We consider the recovery of $u^{\dagger}$ from noisy data $v=Ku+\eta$, where $\eta\in\mathbb{R}^{d}$ is random Gaussian noise. We consider kernels of size $\kappa=2i-1$ for $i=1,\ldots,100$, and random Gaussian noise $\eta\sim\mathcal{N}(0,\sigma^{2}I)$, with $10$ different noise levels, taken at equidistant points in $[10^{-4},10^{-1}]$. Note that for $i=1$ the kernel matrix $K$ is the identity matrix, and as $i$ increases so does the correlation between rows of $K$.
To recover $u^{\dagger}$ from the observation $v$, we minimise $\Phi(x)=g(x)+h(x)$ with $x=u$, $g(x)=\mu{|\kern-1.125pt|}x{|\kern-1.125pt|}_{1}$ and $h(x)=\frac{1}{2}{|\kern-1.125pt|}Kx-v{|\kern-1.125pt|}_{2}^{2}$. The objective $\Phi$ is for SAGA rewritten into the finite-sum form $\Phi(x)=g(x)+\sum_{i=1}^{n}h_{i}(x)$ where $n=d$, $h_{i}(x):=\frac{1}{2}(K_{i}x-v_{i})^{2}$ and $K_{i}\in\mathbb{R}^{1\times n},i=1,...,n$ is the $i$-th row of $K$. Note that $L=1$, ${|\kern-1.125pt|}K_{i}{|\kern-1.125pt|}=1/\sqrt{\kappa},i=1,...,n$ and thus $L_{\max}=1/\kappa$. The regularisation parameter $\mu$ is chosen as $\mu=\frac{1}{2}{|\kern-1.125pt|}K^{*}v{|\kern-1.125pt|}_{\infty}$. We use two step-size choices for both methods. For PGD we use $\tau=1/L$ (denoted PGD1) and $\tau=1.9/L$ (denoted PGD2). For SAGA, we use $\tau=1/(3nL_{\max})$ (denoted SAGA1) and $\tau=1/(1.75nL_{\max})$ (denoted SAGA2).
@@
For each kernel size and noise level we compute a high-precision solution $x^{\ast}$ and then run PGD and SAGA using the stopping criterion ${{|\kern-1.125pt|}x^{(k)}-x^{\ast}{|\kern-1.125pt|}}/{{|\kern-1.125pt|}x^{\ast}{|\kern-1.125pt|}}\leq 10^{-4}.$ We record the number of data passes needed to reach that accuracy, denote it as $N_{\text{PGD}}$ and $N_{\text{SAGA}}$ for the corresponding choices of PGD and SAGA, respectively, and compute the ratio $N_{\text{SAGA}}/N_{\text{PGD}}$, which is averaged over the 10 noise levels. The resulting $4$ ratios (for each step-size pairing), and a line $1/\kappa_{i}$, for $i=1,\ldots,100$ are shown in Figure 1.
@@
We can reach the following observations from the results. First, when $K$ is the identity matrix (for $\kappa=1$, PGD with $\tau=1/L$ converges in one step. Hence, the ratio $N_{\text{SAGA}}/N_{\text{PGD}}$ is very large. Moreover, SAGA1 is faster than SAGA2. This behaviour can be attributed to the properties of $K$. When $\kappa$ is small the gradients $\nabla h_{i}$ and rows of $K$ are less correlated, decreasing the data redundancy. In such a setting stochastic methods can lose some of their competitive edge. On the other side of the coin is the case of large $\kappa$. In this setting the the rows of $K$ are more correlated, which benefits stochastic methods since each $K_{i}$ carries a much larger proportion of the informational content of $K$. In such a setting stochastic methods are expected to perform better. Moreover, larger step-size leads to a larger initial variance and slower performance. Second, for kernel sizes between $5$ and $199$, we observe that the ratio $N_{\text{SAGA}}/N_{\text{PGD}}$ is proportional to $1/\kappa$. Moreover, larger step-size leads to faster performance for both PGD and SAGA. To see this, notice that the orange line $N_{\text{SAGA1}}/N_{\text{PGD2}}$ is above the red line $N_{\text{SAGA1}}/N_{\text{PGD1}}$, indicating that PGD1 needs more data passes to reach stopping criterion. Similarly, the green line stays below the red line, indicating that compared to SAGA2, SAGA1 requires more data passes to reach the stopping criterion.
@@
Figure 1: The lines represent the ratio of the number of data passes for SAGA and PGD, i.e. $N_{\text{SAGA}}/N_{\text{PGD}}$, to reach a target reconstruction accuracy for the deblurring of sparse spikes, for different step-size choices. See accompanying text in Section 5.1 for more details.
5.2   Simulated Parallel-Beam CT
@@
We reproduce here the numerical results from an upcoming publication on a software framework for stochastic optimisation methods for large-scale imaging problems within the Core Imaging Library (CIL) [JAB+21, PAD+21, PJP+24]. This is being developed as a community effort around the CCPi (www.ccpi.ac.uk) and SyneRBI (www.ccpsynerbi.ac.uk) projects. The reproduced results deal with the CT reconstruction of a Shepp–Logan phantom of size $128\times 128$ px. The measurements follow the forward problem $v=Ku^{\dagger}+\eta$, where $\eta$ is a random sample from $\mathcal{N}(0,\sigma^{2}I)$, and the forward operator $K$ is defined by parallel-beam CT with $240$ angles from the $[0,\pi)$ angle range.
@@
We compare the performance of a deterministic algorithm PGD (Algorithm 1) with SGD (Algorithm 7) and SAGA (Algorithm 8). As discussed in Section 3, when solving inverse problems it is not common to utilise fully stochastic approaches, in the sense that instead of using randomly constructed batches the measurement data is pre-partitioned into batches. One element of the partition is then sampled at random in each iteration.
@@
To recovert the phantom $u^{\dagger}$ we minimise $\Phi(x)=g(x)+h(x)$, where $x=u$, $g(x)=\lambda\text{TV}(x)$, and $h(x)=\frac{1}{2}\|Kx-v\|^{2}$ for PGD. SGD and SAGA rewrite the objective as $\Phi(x)=g(x)+\sum_{i=1}^{n}h_{i}(x)$, where $n\in\mathbb{N}$, and $h_{i}(x)=\frac{1}{2}\|K_{i}x-v_{i}\|^{2}$ is such that $\sum_{i=1}^{n}h_{i}(x)=h(x)$. Operator $K_{i}$, for $1\leq i\leq n$, is defined as the CT operator that uses $240/n$ angles from the range $[(i-1)/n,\pi)$ by taking every $n$-th angle starting with $i/n$ with increments of size $\pi n/240$. This is sometimes called the equidistant or staggered partition, and results in all the operator norms $\|K_{i}\|$ to be of a nearly identical value. Moreover, the number of batches is typically a divisor of the number of angles, which ensures that all the batches are of the same size.
@@
(c) Reconstructions after 10 data passes with PGD, SGD and SAGA
@@
In Figure 2 we show the performance of these 3 algorithms in terms of the sub-optimality of the objective value: $\Phi(x^{(k)})-\Phi(x^{\ast})$, where $\Phi(x^{\ast})$ is a reference solution computed with FISTA (Algorithm 2). Figure 2 shows a significant speed up with SGD and SAGA over PGD. SGD is the fastest method in early iterations, but it then slows down after $5$-$10$ data passes. On the other hand, SAGA is steady and fast throughout the optimisation trajectory and the increased variance it presents in initial iterations disappears already after around $20$ data passes. Figure 2a and Figure 2b also investigate the effect of the number of subsets: the lightest shade shows a low number of subsets ($n=10$), normal shade shows a higher number of subsets ($n=60$), and the darkest shade shows the highest number of subsets ($n=240$). We observe that increasing the number of subsets has a dual effect: it increases the variance of the optimisation trajectory for stochastic algorithms, but it greatly increases the speed. In Figure 2c we show the reconstructions of the 4 studied algorithms obtained after $10$ data passes. The reconstruction with PGD presents as oversmooth, and fails to adequately reconstruct the finer image detail. For example, consider the zoomed in region where stochastic algorithms manage to recover two distinct objects (one of a lower and the other of a higher intensity), whereas PGD does not.
@@
Figure 3: Comparison of the performance of stochastic and incremental methods (using the Herman–Meyer order) on parallel-beam CT reconstruction of the Shepp-Logan
In Figure 3 we compare the effect of the strategy for accessing the function indices. Specifically, we compare SGD and SAGA, with indices sampled uniformly at random (with replacement), and IGD and IAGA, which are incremental gradient formulations of SGD and SAGA that access indices according to a deterministic Herman–Meyer order [HM93]. The results show that the behaviour highly depends on the number $n$ of subsets and the used algorithm. Comparing SGD and IGD reveals that the deterministic strategy can outperform the stochastic strategy, though the differences are not significant. However, for variance reduced algorithms the differences are significantly more severe. When $n$ is small, SAGA outperforms IAGA and presents a much smaller variance, though the variance with IAGA reduces significantly within the first $50$ data passes. When $n$ is large, SAGA is significantly better than IAGA.
5.3   Fan-Beam CT Data of a Walnut
In this set of experiments we take the central slice of a cone beam CT scan of a walnut [Mea20]. As in Section 5.2, we reproduce the numerical results from an upcoming publication on the stochastic framework in the CIL software library. Full sinogram data has 721 projections in the $[0,2\pi)$ range. To reduce the computational costs we consider only $360$ projections in the $[0,2\pi)$ range, with a $1^{\circ}$ angle separation. The reconstructed images are of size $280\times 280$ px.
Noisy measurements are generated by corrupting the sinogram (the measurement data) following the Beer–Lamberts law. We first compute the expected number of counts via $N_{I}=\text{Pois}(I_{0}\exp(-v))$, where $v$ is the loaded (noiseless) sinogram and $I_{0}$ the intensity of the initial beam. The noisy measurement is then computed by first correcting all 00 entries in $N_{I}$ to $1$, and then applying linearisation (post-log measurements) as $v^{\delta}=-\ln(N_{I}/I_{0})$. We consider three intensity levels: low intensity ($I_{0}=50$), medium intensity ($I_{0}=250$), and high intensity ($I_{0}=5000$), simulating high to low noise. The objective is defined as